OpenStudy (anonymous):

what would the slope be in -2x+3y=15

OpenStudy (anonymous):

To find the slope you'll want to solve the equation for y, so you have y by itself on one side of the equation

OpenStudy (anonymous):

exactly...put your equation in y = mx + b form where m is your slope

OpenStudy (anonymous):

you do noy need to solve the equation to find the slope

OpenStudy (anonymous):

typo...not

OpenStudy (anonymous):

3y=15+2x divide both sides by 3 y=15/3+2x/3 y=5+2/3X slope is 2/3

OpenStudy (anonymous):

^thank you for breaking it down and explaining it to me

OpenStudy (anonymous):

:)